OpenStudy (anonymous):

(-3,-5), parallel to y=-4x+3

OpenStudy (anonymous):

point-slope form: (y-k) = m (x-h) where (h,k) is your point and m is the slope the slope of the line y=-4x+3 is (-4) so, (y - (-5)) = (-4) (x - (-3)) y + 5 = (-4) (x + 3) ...